# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### v5

@@ 5 More Technical Details
Specifically, we describe how we calculate the Global Significance Score for each trained Gaussian in Algorithm 2.
@@
6 Detailed Experiment Results
@@
Tables 7, 8, and 9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 724MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
Figure 8: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [40]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Figure 9: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [40]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Table 7: Quantitative Comparison (PSNR) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
@@
3D-GS 25.122 21.324 27.294 26.783 31.687 22.500 29.114 31.618 32.307 24.778 21.449
Ours 25.093 21.252 27.043 26.862 31.680 22.658 28.464 30.974 31.459 24.851 21.407
Implementation Details of VQ-DVGO.
Table 8: Quantitative Comparison (SSIM) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
@@
3D-GS 0.746 0.574 0.856 0.770 0.926 0.630 0.914 0.932 0.946 0.863 0.781
Ours 0.745 0.566 0.845 0.775 0.923 0.627 0.900 0.921 0.938 0.862 0.772
Table 9: Quantitative Comparison (LPIPS) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
@@
3D-GS 0.245 0.381 0.122 0.242 0.200 0.366 0.186 0.118 0.184 0.175 0.262
Ours 0.257 0.392 0.143 0.251 0.212 0.378 0.211 0.140 0.199 0.179 0.283
7 Conclusion and Future Works
We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis. For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification. In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer. A 3D Gaussian Vector Quantization post-processing effectively removes further redundancy by using Gaussian significance as an indicator. With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss. Exploring the possibility of zero-shot compression across various 3D-GS-based frameworks is a valuable direction for future research.

### v6

@@ 5 Conclusion, Limitations, and Broad Impact
Specifically, we detail how we calculate each Gaussian’s Global Significance Score in Algorithm 2.
@@
6 More Experiment Results
@@
Figure 7: Visual Comparisons for Ablation Study. We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab. The final model (Exp #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss. To highlight the difference, we visualize the SSIM between the rendered images and the GT, where a whiter output in the residual signifies closer alignment with the GT, denoting better quality, and the area with greater color intensity indicates lower quality.
@@
Tables 8, 9, and 10 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 782MB to 45MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
Figure 8: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [16], LightGaussian, and VQ-DVGO [49]. The residual maps highlight the differences between the rendered images and ground truth (GT) images.
Figure 9: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [16], LightGaussian, and VQ-DVGO [49]. The residual maps highlight the differences between the rendered images and ground truth (GT) images.
Table 8: Quantitative Comparison (PSNR) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS, Compressed 3D-GS, Compact 3D-GS using their original code and 3D-GS configurations on our platform, to perform a fair comparison.
@@
3D-GS 25.194 21.414 27.29 26.598 31.336 22.476 28.989 31.328 31.948 25.372 21.957
Compressed 3D-GS 25.040 21.126 26.817 26.357 31.072 22.298 28.653 30.755 31.161 25.194 21.882
Compact 3D-GS 24.795 20.974 26.703 26.260 30.645 22.570 28.553 30.332 31.84 25.041 21.544
Ours 25.196 21.538 26.961 26.77 31.399 22.685 28.478 30.869 31.414 25.399 21.838
Table 9: Quantitative Comparison (SSIM) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS, Compressed 3D-GS, Compact 3D-GS using their original code and 3D-GS configurations on our platform, to perform a fair comparison.
@@
3D-GS 0.764 0.601 0.862 0.770 0.917 0.632 0.906 0.925 0.939 0.878 0.811
Compressed 3D-GS 0.752 0.584 0.845 0.757 0.912 0.622 0.896 0.916 0.932 0.873 0.803
Compact 3D-GS 0.737 0.568 0.837 0.753 0.909 0.632 0.892 0.913 0.931 0.869 0.788
Ours 0.765 0.598 0.855 0.776 0.916 0.635 0.898 0.919 0.934 0.877 0.801
Table 10: Quantitative Comparison (LPIPS) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS, Compressed 3D-GS, Compact 3D-GS using their original code and 3D-GS configurations on our platform, to perform a fair comparison.
@@
3D-GS 0.211 0.339 0.108 0.216 0.221 0.327 0.201 0.127 0.206 0.147 0.209
Compressed 3D-GS 0.237 0.358 0.139 0.248 0.234 0.352 0.215 0.140 0.219 0.157 0.220
Compact 3D-GS 0.256 0.378 0.144 0.256 0.233 0.347 0.223 0.140 0.218 0.163 0.240
Ours 0.218 0.352 0.122 0.222 0.232 0.338 0.220 0.141 0.221 0.155 0.239
Implementation Details of VQ-DVGO.
Overall Analysis.